# OPENXRD: A Comprehensive Benchmark and Enhancement Framework for LLM/MLLM XRD Question Answering

### v1

@@ 3.3 Open-Book Mode Observations Using AI-Generated Materials (No Expert Review)
Table 2 lists the closed-book mode vs. open-book mode accuracy on our 217-question crystallography benchmark. Here and in subsequent tables, $\Delta$ represents the performance improvement defined in Equation (4), with model subscripts omitted as each row corresponds to a different model. We first evaluate open-book mode accuracy by giving each model relevant supporting textual material generated by GPT-4.5. The largest improvements occur in smaller LLaVA-based models, which gain 4–6% from the supporting textual material clarifying domain specifics. O3-mini sees only a minor jump (+0.46%), possibly because it already "knows" the domain. GPT-4.5-preview slightly decreases in open-book mode (-2.77%), suggesting occasional confusion from extraneous supporting textual material or contradictory references. Figure 6 shows an example where supporting materials have caused confusion. In this case, a model with correct closed-book knowledge has been misled by materials mentioning several related but distinct concepts. Our detailed analysis of the AI-generated supporting materials revealed several limitations: occasional technical inaccuracies in specialized crystallographic terminology, inadequate depth of explanation for particularly complex phenomena, information presented at times being too general to help with specific question nuances, and in some cases, the supporting materials were tangential to the precise knowledge needed for the question. These observations led us to hypothesize that expert-reviewed materials might yield greater performance improvements, particularly for mid-sized models with significant capacity but incomplete domain knowledge.
Question Which scattering mechanism causes damping of the atomic scattering factor at higher angles? Options
(A) Compton scattering
(B) Electron cloud interference
(C) Thermal vibration
(D) Core-level absorption
Model Responses LLaVA-v1.6-34B (Closed-book): B ✓ LLaVA-v1.6-34B (Open-book): C ✗ Supporting Material Excerpt “…The atomic scattering factor decreases with increasing scattering angle due to interference effects within the electron cloud. Additionally, thermal vibrations (Debye-Waller factor) can further dampen the intensity at higher angles, though this is a separate effect from the intrinsic angular dependence…”
@@
Fig.  6: Example of confusion in LLaVA-v1.6-34B due to conflicting supporting material. The model had correct knowledge in closed-book mode but was misled by additional content that mentioned both the correct answer and a distractor as possible factors affecting scattering, albeit in different ways.
@@
Table 2: Comparison of model accuracy in closed-book mode vs. open-book mode with AI-generated supporting materials. $\Delta=$ (Open-book mode) -- (Closed-book mode).
@@
GPT-4.5-preview 93.09 90.32 -2.77
O3-mini 88.94 89.40 +0.46
O1 87.56 88.02 +0.46
@@
GPT-4 81.11 82.03 +0.92
LLaVA-v1.6-34B 66.80 72.81 +4.61
LLaVA-onevision-QWEN2-7B-si 66.40 71.89 +5.53
LLaVA-onevision-QWEN2-7B-ov-chat 65.90 72.35 +6.45
LLaVA-onevision-QWEN2-7B-ov 65.40 71.43 +5.99
LLaVA-v1.6-mistral-7B 53.00 58.53 +5.75
LLaVA-onevision-QWEN2-0.5B-ov 47.50 51.15 +3.68
LLaVA-v1.5-13B 46.50 44.24 -2.30
LLaVA-onevision-QWEN2-0.5B-si 46.10 51.15 +5.07
LLaVA-v1.5-7B 18.00 23.96 +5.99
@@ 3.4 Open-Book Mode Observations Using AI-Generated Materials with Expert Review
Building on the findings from the initial open-book evaluations, we introduced expert review to improve the quality and effectiveness of the supporting materials. Three PhD-level crystallography experts were engaged to refine the AI-generated materials. Table 3 compares model accuracy in closed-book and open-book modes using expert-reviewed supporting materials across our 217-question crystallography benchmark. The results show substantially improved performance, particularly among LLaVA-based models.
The largest accuracy gains are observed in LLaVA-v1.6-34B (+11.50%) and LLaVA-v1.6-mistral-7B (+11.10%), highlighting the effectiveness of expert-curated guidance in boosting model reasoning. Several other mid-sized models, such as LLaVA-onevision-QWEN2-7B-ov-chat (+10.10%) and LLaVA-onevision-QWEN2-7B-ov (+8.80%), also demonstrate notable improvements.
Interestingly, GPT-4.5-preview exhibits a slight performance drop (-3.69%) even with expert-reviewed texts. This suggests that highly capable models may already possess sufficient internal knowledge or may be more sensitive to redundant or overly detailed inputs.
@@
Table 3: Comparison of model accuracy in closed-book mode vs. open-book mode with expert-reviewed supporting materials. $\Delta=$ (Open-book mode) -- (Closed-book mode).
@@
GPT-4.5-preview 93.09 89.40 -3.69
O3-mini 88.94 89.90 +0.96
O1 87.56 88.90 +1.34
GPT-4-turbo-preview 83.41 84.80 +1.39
GPT-4 81.11 82.50 +1.39
LLaVA-v1.6-34B 66.80 78.30 +11.50
LLaVA-onevision-QWEN2-7B-ov-chat 65.90 76.00 +10.10
LLaVA-onevision-QWEN2-7B-si 66.40 75.10 +8.70
LLaVA-onevision-QWEN2-7B-ov 65.40 74.20 +8.80
LLaVA-v1.6-mistral-7B 53.00 64.10 +11.10
LLaVA-onevision-QWEN2-0.5B-ov 47.50 54.40 +6.90
LLaVA-v1.5-13B 46.50 48.40 +1.90
LLaVA-onevision-QWEN2-0.5B-si 46.10 53.90 +7.80
LLaVA-v1.5-7B 18.00 28.60 +10.60
3.5 Comparative Analysis of AI-Generated Supporting Materials: Without vs. With Expert Review
To better understand the impact of expert review, Table 4 presents a direct comparison of performance gains ($\Delta$) between AI-generated supporting materials without expert review and those refined through expert review. Our comparative analysis reveals several significant patterns and Key Findings from Comparative Analysis. A near-universal improvement was observed with expert review, where 13 out of 14 models showed greater performance gains or smaller losses compared to those using AI-generated materials. Specifically, mid-capacity models like LLaVA-v1.6-34B and LLaVA-v1.6-mistral-7B saw the most dramatic improvements, with additional gains of 6.89% and 5.35% respectively. Conversely, models such as LLaVA-v1.5-13B and GPT-4-turbo-preview, which initially showed negative effects with AI-generated materials, experienced positive effects with expert-reviewed ones. However, top models already near ceiling performance, such as GPT-4.5-preview, O3-mini, and O1, saw minimal additional benefits or even slight degradation. Performance differences suggest that expert-reviewed materials successfully addressed the limitations identified in Phase 1, particularly for models with sufficient capacity to utilize improved information but insufficient internal knowledge to solve problems independently. Furthermore, Figure 7 illustrates a categorized performance comparison of these two phases of simulations.

### v2

@@ 3.3 Open-Book Mode Observations Using AI-Generated Materials (No Expert Review)
Table 3 lists the closed-book mode vs. open-book mode accuracy on our 217-question crystallography benchmark. Here and in subsequent tables, $\Delta$ represents the performance improvement defined in Equation (4), with model subscripts omitted as each row corresponds to a different model. We first evaluate open-book mode accuracy by giving each model relevant supporting textual material generated by GPT-4.5.
We observe distinct performance patterns across model scales, architectures, and domain specializations. Remarkably, specialized coder models demonstrate the most dramatic improvements: qwen3-coder-plus gains +65.95% and qwen3-coder-flash gains +40.87% from AI-generated supporting materials, suggesting these models have significant reasoning capacity but minimal crystallographic knowledge. Among general-purpose models, mid-sized architectures (>7B–70B parameters) show substantial benefits, with models like llama-3.1-8b-instruct (+6.00%), llava-onevision-qwen2-7b-ov-chat (+6.45%), anthropic/claude-3-haiku (+8.45%), and llava-v1.6-34b (+5.99%) gaining meaningfully from domain-specific context. Larger models (70B+) such as llama-3.1-405b-instruct (-4.61%), llama-3-70b-instruct (-4.14%), and qwen-2.5-72b-instruct (-1.38%) show minimal improvement or slight degradation, while very large frontier models like o3-mini see only minor gains (-4.15%), suggesting they already possess substantial internal crystallographic knowledge.
Notably, multiple frontier models show performance degradation with AI-generated supporting materials: GPT-4.5-preview (-2.31%), Claude-3.5-Sonnet (-1.84%), dziner-qwen-2.5-72b (-3.22%), and mistral-large (-2.31%). This counterintuitive pattern reveals an important failure mode: for models with comprehensive pre-trained knowledge, additional context—even when domain-relevant—can introduce interference rather than assistance.
Figure 8 shows a concrete example where supporting materials have caused confusion, directly illustrating this interference mechanism. In this case, LLaVA-v1.6-34B answered correctly in closed-book mode (Option B: Electron cloud interference) but was misled in open-book mode to select Option C (Thermal vibration). The supporting material discussed both phenomena in close proximity: “The atomic scattering factor decreases with increasing scattering angle due to interference effects within the electron cloud. Additionally, thermal vibrations (Debye-Waller factor) can further dampen the intensity…” This juxtaposition caused the model to conflate two distinct mechanisms—the intrinsic angular dependence (correct answer) versus temperature-dependent effects (distractor)—demonstrating that the failure mode is primarily conceptual interference and distraction from related concepts rather than simple information overload or redundancy.
Our detailed analysis of the AI-generated supporting materials revealed several limitations: occasional technical inaccuracies in specialized crystallographic terminology, inadequate depth of explanation for particularly complex phenomena, information presented at times being too general to help with specific question nuances, and in some cases, the supporting materials were tangential to the precise knowledge needed for the question. Figure 7 summarizes these patterns by model size group, showing that mid-sized and small models benefit substantially from AI-generated materials ($5.3\%$ and $10.3\%$ respectively), while large models experience slight degradation ($1.9\%$).
These observations led us to hypothesize that expert-reviewed materials might yield greater performance improvements, particularly for mid-sized models with significant capacity but incomplete domain knowledge.
Fig. 7: Open-book performance with AI-generated supporting materials by model size group. Large Models (>>70B parameters or advanced architectures, n=27); Mid-sized Models (>>7B-70B parameters, n=22); Small Models ($\leq$7B parameters, n=25). Compared to closed-book performance (Figure 6), large models show slight degradation ($1.9\%$), while mid-sized models gain $5.3\%$ and small models gain $10.3\%$, demonstrating that AI-generated context benefits models with knowledge gaps but can interfere with already-capable models.
@@
Fig. 8: Example of confusion in LLaVA-v1.6-34B due to conflicting supporting material. The model had correct knowledge in closed-book mode but was misled by additional content that mentioned both the correct answer and a distractor as possible factors affecting scattering, albeit in different ways.
@@
Table 3: Comparison of model accuracy in closed-book mode vs. open-book mode with AI-generated supporting materials. $\Delta=$ (Open-book mode) - (Closed-book mode).
@@
openai/gpt-5 96.77 94.93 -1.84
x-ai/grok-4-fast 96.31 95.39 -0.92
openai/gpt-5-codex 96.31 94.47 -1.84
google/gemini-2.5-pro 95.39 88.48 -6.91
o3-mini 93.55 89.40 -4.15
meituan/longcat-flash-chat 93.09 90.32 -2.77
gpt-4.5-preview 92.63 90.32 -2.31
anthropic/claude-3.5-sonnet 91.24 89.40 -1.84
openai/gpt-4o 90.74 87.56 -3.18
dziner-qwen-2.5-72b 90.32 87.10 -3.22
perplexity/sonar-pro 90.32 90.32 +0.00
qwen/qwen3-next-80b-a3b-instruct 90.32 89.40 -0.92
deepseek/deepseek-v3.1-terminus 89.86 90.78 +0.92
qwen/qwen-plus 89.86 90.78 +0.92
qwen/qwen3-next-80b-a3b-thinking 89.86 87.10 -2.76
deepseek/deepseek-chat 89.40 91.24 +1.84
o1 89.40 88.02 -1.38
qwen/qwen-max 89.40 88.02 -1.38
anthropic/claude-3-opus 88.94 88.02 -0.92
openai/o1-mini 88.02 88.48 +0.46
meta-llama/llama-3.1-405b-instruct 87.56 82.95 -4.61
qwen/qwen-2.5-72b-instruct 87.56 86.18 -1.38
mistralai/mistral-large 86.18 83.87 -2.31
amazon/nova-pro-v1 86.11 86.64 +0.53
gpt-4-turbo 85.25 85.71 +0.46
meta-llama/llama-3-70b-instruct 84.79 80.65 -4.14
meta-llama/llama-3.1-70b-instruct 84.79 83.41 -1.38
dziner-qwen-2.5-coder-32b 83.87 84.79 +0.92
openai/gpt-4-0314 83.41 77.42 -5.99
@@
amazon/nova-lite-v1 82.03 84.79 +2.76
dziner-qwen-2.5-7b 81.57 79.26 -2.31
gpt-4 81.57 82.03 +0.46
openai/gpt-4o-mini 81.11 82.03 +0.92
openrouter/auto 81.11 81.57 +0.46
qwen/qwen-2.5-7b-instruct 79.72 75.12 -4.60
google/gemma-2-27b-it 79.26 78.80 -0.46
anthropic/claude-3.5-haiku 77.57 88.02 +10.45
mistralai/mixtral-8x22b-instruct 76.81 80.09 +3.28
mistralai/mistral-7b-instruct 75.35 77.42 +2.07
google/gemma-2-9b-it 75.12 77.42 +2.30
amazon/nova-micro-v1 74.65 79.26 +4.61
mistralai/mistral-small 74.65 79.26 +4.61
anthropic/claude-3-haiku 74.04 82.49 +8.45
openai/gpt-3.5-turbo-16k 72.60 71.01 -1.59
mistralai/mixtral-8x7b-instruct 72.33 75.96 +3.63
meta-llama/llama-3-8b-instruct 71.89 73.27 +1.38
mistralai/pixtral-12b 71.83 75.58 +3.75
openai/gpt-3.5-turbo 70.67 71.50 +0.83
meta-llama/llama-3.1-8b-instruct 69.12 75.12 +6.00
llava-v1.6-34b 66.82 72.81 +5.99
lmms-lab/llava-onevision-qwen2-7b-si 66.36 71.89 +5.53
lmms-lab/llava-onevision-qwen2-7b-ov-chat 65.90 72.35 +6.45
lmms-lab/llava-onevision-qwen2-7b-ov 65.44 71.43 +5.99
meta-llama/llama-3.2-3b-instruct 64.98 62.67 -2.31
arcee-ai/afm-4.5b 62.21 64.52 +2.31
mistralai/mistral-7b-instruct-v0.1 59.50 54.03 -5.47
perplexity/sonar 59.24 68.84 +9.60
undi95/remm-slerp-l2-13b 57.08 61.50 +4.42
llamat-3-chat 57.14 30.41 -26.73
alibaba/tongyi-deepresearch-30b-a3b 55.76 80.18 +24.42
gryphe/mythomax-l2-13b 53.77 61.68 +7.91
llava-v1.6-mistral-7b 52.99 58.53 +5.54
llamat-2-chat 50.69 21.13 -29.56
lmms-lab/llava-onevision-qwen2-0.5b-ov 47.47 51.15 +3.68
llava-v1.5-13b 46.54 44.24 -2.30
lmms-lab/llava-onevision-qwen2-0.5b-si 46.08 51.15 +5.07
qwen/qwen3-coder-flash 43.46 84.33 +40.87
qwen/qwen3-coder-plus 22.99 88.94 +65.95
honeybee-13b 22.12 23.04 +0.92
qwen/qwen-2.5-coder-32b-instruct 21.03 27.01 +5.98
honeybee-7b 19.35 23.04 +3.69
llava-v1.5-7b 17.97 23.96 +5.99
llamat-2 16.59 19.82 +3.23
@@ 3.4 Open-Book Mode Observations Using AI-Generated Materials with Expert Review
Building on the findings from the initial open-book evaluations, we introduced expert review to improve the quality and effectiveness of the supporting materials. Three PhD-level crystallography experts were engaged to refine the AI-generated materials.
Table 4 compares model accuracy in closed-book and open-book modes using expert-reviewed supporting materials across our 217-question crystallography benchmark. The results show substantially improved performance, particularly among LLaVA-based models.
Remarkably, specialized coder models show the most dramatic gains, with qwen3-coder-plus (+65.95%) and qwen3-coder-flash (+42.25%) achieving exceptional improvements from very low baselines. Among more general-purpose vision-language models, the largest accuracy gains are observed in LLaVA-v1.6-34B (+11.52%) and LLaVA-v1.6-mistral-7B (+11.07%), highlighting the effectiveness of expert-curated guidance in boosting model reasoning.
Several other mid-sized models, such as LLaVA-onevision-QWEN2-7B-ov-chat (+10.14%) and LLaVA-onevision-QWEN2-7B-ov (+8.75%), also demonstrate notable improvements. Figure 9 illustrates these gains aggregated by model size group, showing that expert curation provides additional improvements over AI-generated materials, particularly for mid-sized ($1.7\%$) and small models ($2.5\%$), while large models maintain stable performance.
A notable failure mode identified in our analysis is catastrophic degradation in domain-specialized models, which experience substantial performance drops despite receiving expert-reviewed supporting materials. As summarized in Tables 34, the LLaMAT family—pretrained on materials-science literature and domain-focused instruction datasets—shows the most severe declines. LLaMAT-3-chat falls from 57.14% (closed-book) to 31.80% (open-book with expert review), a –25.34% drop, while LLaMAT-2-chat decreases from 50.69% to 16.13%, a –34.56% degradation that renders the model nearly unusable in open-book mode. Only the base llamat-2 model shows a modest improvement (+3.23%), though its low absolute accuracy indicates fundamental capacity limitations regardless of external support. These results demonstrate that domain specialization alone does not guarantee effective external-knowledge integration.
Similar patterns emerge in other domain-focused families. The HoneyBee models, designed primarily for scientific figure interpretation, show limited transferability to text-only crystallographic reasoning. HoneyBee-13B drops from 22.12% to 21.20% (–0.92%), and HoneyBee-7B from 19.35% to 17.51% (–1.84%), suggesting that pretraining on visual-scientific corpora does not enhance performance on purely textual tasks, and that their small parameter scale (7B–13B) restricts knowledge retention. The dZiner family, consisting of QWEN variants fine-tuned on materials-science corpora, exhibits a similar sensitivity to interference. While dZiner-QWEN-2.5-72B attains strong closed-book accuracy (90.32%), it drops to 86.18% (–4.14%) with expert-reviewed materials. The medium-sized dZiner-QWEN-2.5-coder-32B declines by –6.45%, and the smaller dZiner-QWEN-2.5-7B by –2.31%. These negative open-book effects mirror degradation trends observed in other high-capacity models, indicating that domain-specific pretraining raises baseline performance but simultaneously increases vulnerability to information interference.
Across all three families, the underlying mechanism is representational rigidity introduced during domain adaptation. These models internalize crystallographic and materials-science concepts in narrowly defined textual, visual, or instructional formats. When exposed to expert-reviewed passages that express these concepts using more pedagogical, descriptive, or stylistically diverse language, they encounter distributional mismatch between their internal priors and the external context. Even when the supplemental information is factually correct, this mismatch triggers interference rather than assistance, leading to performance degradation. This reveals a fundamental trade-off in domain specialization: while narrow, domain-focused pretraining can strengthen closed-book accuracy, it reduces flexibility in assimilating new or differently framed knowledge sources.
Fig. 9: Open-book performance with expert-reviewed supporting materials by model size group. Large Models (>70B parameters or advanced architectures, n=27); Mid-sized Models (>7B-70B parameters, n=22); Small Models ($\leq$7B parameters, n=25). Compared to AI-generated materials (Figure 7), expert curation provides additional gains for mid-sized ($1.7\%$) and small models ($2.5\%$), while large models remain stable, demonstrating that content quality improvements benefit models with reasoning capacity but incomplete domain knowledge.
Interestingly, the knowledge-interference failure pattern extends systematically across top-tier models, as explained in Section 3.3. As summarized in Table 4, in the transition from the closed-book setting to the open-book mode with expert-reviewed supporting materials, several frontier systems show measurable degradation despite their high baseline accuracy. GPT-4.5-preview declines from 92.63% to 89.40% (–3.23 percentage points), while other state-of-the-art models exhibit comparable or greater decreases: openai/gpt-5 (–3.68%), x-ai/grok-4-fast (–5.78%), openai/gpt-5-codex (–4.14%), and O3-mini (–3.69%). Importantly, all of these systems exceed 90% closed-book accuracy, confirming that they already possess extensive crystallographic knowledge. The consistency of this degradation across diverse architectures (spanning distinct training paradigms, model families, and organizations) indicates that knowledge interference is an intrinsic characteristic of high-capacity language models rather than a model-specific artifact. These models appear highly sensitive to redundant or overlapping external information: when provided with additional context that reformulates or reiterates knowledge already internalized, they can misallocate attention, resulting in subtle yet systematic declines in performance even when the supplemental material is accurate and expert-curated.
@@
Table 4: Comparison of model accuracy in closed-book mode vs. open-book mode with expert-reviewed supporting materials. $\Delta=$ (Open-book mode) - (Closed-book mode).
@@
openai/gpt-5 96.77 93.09 -3.68
x-ai/grok-4-fast 96.31 90.53 -5.78
openai/gpt-5-codex 96.31 92.17 -4.14
google/gemini-2.5-pro 95.39 92.63 -2.76
o3-mini 93.55 89.86 -3.69
meituan/longcat-flash-chat 93.09 89.86 -3.23
gpt-4.5-preview 92.63 89.40 -3.23
anthropic/claude-3.5-sonnet 91.24 89.40 -1.84
openai/gpt-4o 90.74 87.10 -3.64
dziner-qwen-2.5-72b 90.32 86.18 -4.14
qwen/qwen3-next-80b-a3b-instruct 90.32 89.40 -0.92
perplexity/sonar-pro 90.32 88.02 -2.30
deepseek/deepseek-v3.1-terminus 89.86 90.78 +0.92
qwen/qwen-plus 89.86 89.40 -0.46
qwen/qwen3-next-80b-a3b-thinking 89.86 87.10 -2.76
deepseek/deepseek-chat 89.40 90.32 +0.92
qwen/qwen-max 89.40 89.40 +0.00
o1 89.40 88.94 -0.46
anthropic/claude-3-opus 88.94 88.94 +0.00
openai/o1-mini 88.02 88.02 +0.00
meta-llama/llama-3.1-405b-instruct 87.56 84.33 -3.23
qwen/qwen-2.5-72b-instruct 87.56 85.71 -1.85
mistralai/mistral-large 86.18 86.18 +0.00
amazon/nova-pro-v1 86.11 86.64 +0.53
gpt-4-turbo 85.25 85.71 +0.46
meta-llama/llama-3-70b-instruct 84.79 82.95 -1.84
meta-llama/llama-3.1-70b-instruct 84.79 85.19 +0.40
dziner-qwen-2.5-coder-32b 83.87 77.42 -6.45
openai/gpt-4-0314 83.41 82.49 -0.92
gpt-4-turbo-preview 83.41 84.79 +1.38
amazon/nova-lite-v1 82.03 85.71 +3.68
dziner-qwen-2.5-7b 81.57 79.26 -2.31
gpt-4 81.57 82.49 +0.92
openai/gpt-4o-mini 81.11 81.11 +0.00
openrouter/auto 81.11 81.57 +0.46
qwen/qwen-2.5-7b-instruct 79.72 78.34 -1.38
google/gemma-2-27b-it 79.26 79.72 +0.46
anthropic/claude-3.5-haiku 77.57 85.92 +8.35
mistralai/mixtral-8x22b-instruct 76.81 83.89 +7.08
mistralai/mistral-7b-instruct 75.35 80.65 +5.30
google/gemma-2-9b-it 75.12 81.57 +6.45
amazon/nova-micro-v1 74.65 80.65 +6.00
mistralai/mistral-small 74.65 83.41 +8.76
anthropic/claude-3-haiku 74.04 83.33 +9.29
openai/gpt-3.5-turbo-16k 72.60 75.85 +3.25
mistralai/mixtral-8x7b-instruct 72.33 77.14 +4.81
meta-llama/llama-3-8b-instruct 71.89 75.58 +3.69
mistralai/pixtral-12b 71.83 77.88 +6.05
openai/gpt-3.5-turbo 70.67 76.33 +5.66
meta-llama/llama-3.1-8b-instruct 69.12 78.34 +9.22
llava-v1.6-34b 66.82 78.34 +11.52
lmms-lab/llava-onevision-qwen2-7b-si 66.36 75.12 +8.76
lmms-lab/llava-onevision-qwen2-7b-ov-chat 65.90 76.04 +10.14
lmms-lab/llava-onevision-qwen2-7b-ov 65.44 74.19 +8.75
meta-llama/llama-3.2-3b-instruct 64.98 65.44 +0.46
arcee-ai/afm-4.5b 62.21 67.28 +5.07
mistralai/mistral-7b-instruct-v0.1 59.50 62.44 +2.94
perplexity/sonar 59.24 67.14 +7.90
undi95/remm-slerp-l2-13b 57.08 65.12 +8.04
llamat-3-chat 57.14 31.80 -25.34
alibaba/tongyi-deepresearch-30b-a3b 55.76 86.64 +30.88
gryphe/mythomax-l2-13b 53.77 66.51 +12.74
llava-v1.6-mistral-7b 52.99 64.06 +11.07
llamat-2-chat 50.69 16.13 -34.56
lmms-lab/llava-onevision-qwen2-0.5b-ov 47.47 54.38 +6.91
llava-v1.5-13b 46.54 48.39 +1.85
lmms-lab/llava-onevision-qwen2-0.5b-si 46.08 53.91 +7.83
qwen/qwen3-coder-flash 43.46 85.71 +42.25
qwen/qwen3-coder-plus 22.99 88.94 +65.95
honeybee-13b 22.12 21.20 -0.92
qwen/qwen-2.5-coder-32b-instruct 21.03 30.77 +9.74
honeybee-7b 19.35 17.51 -1.84
llava-v1.5-7b 17.97 28.57 +10.60
llamat-2 16.59 19.82 +3.23
3.5 Ablation Study: Content Quality vs. Information Quantity
To rigorously test whether performance gains stem from information quantity or expert curation quality, we conducted a comprehensive token-matched comparison between AI-generated and expert-reviewed supporting materials. Tokens are the fundamental units of text that language models process—roughly corresponding to words or sub-word pieces (e.g., “crystallography” might be split into “crystal” + “lography”). By controlling token count, we ensure that both material types contain equivalent amounts of text, allowing us to isolate content quality effects from simple volume differences. This ablation directly addresses the question: do expert improvements reflect better content or simply more text?
3.5.1 Token-Matched Experimental Design
To isolate the effects of content quality from those of information quantity, we implemented a token-matched ablation study in which every AI-generated supporting passage was paired with its expert-reviewed counterpart of nearly identical length. This design ensured that both sets of materials provided the same textual volume, allowing any observed performance differences to be directly attributed to qualitative factors such as accuracy, relevance, and pedagogical clarity.
As summarized in Table 5, the mean token count was $566.0\pm 79.2$ for AI-generated materials and $568.9\pm 82.7$ for expert-reviewed materials, representing a negligible difference of only 2.9 tokens (0.51% change). Moreover, the high correlation ($r=0.948$) between AI- and expert-reviewed helper lengths across all 217 questions, along with their nearly identical median and interquartile ranges ($558.0$ [508.0508.0605.0605.0] vs. $558.0$ [513.0513.0610.0610.0]), confirms that the two distributions are tightly matched.
This strict token control eliminates text volume as a potential confounding factor and establishes a high-fidelity basis for quality-driven comparison. In other words, any measurable performance improvements can be confidently attributed to the qualitative enhancements introduced through expert curation, rather than to variations in the amount of information provided. This methodological precision is critical for disentangling the true impact of expert refinement from superficial gains that might otherwise arise from longer or more verbose passages.
Table 5: Token Count Statistics for Questions and Helper Texts. Expert-reviewed materials maintain virtually identical length to AI-generated materials (mean difference: 2.9 tokens, 0.51%), enabling quality-controlled comparison.
Metric Questions AI-Generated Expert-Reviewed
Mean $\pm$ SD $50.1\pm 15.2$ $566.0\pm 79.2$ $568.9\pm 82.7$
Median [IQR] $49.0$ [38.0–60.0] $558.0$ [508.0–605.0] $558.0$ [513.0–610.0]
Range [21–98] [378–919] [378–919]
Total 10,882 122,825 123,457
Difference +2.9 tokens (+0.51%)
Correlation $r=0.948$
3.5.2 LLM-Based Helper Quality Scoring
To quantitatively validate the qualitative advantages of expert-reviewed helpers, we scored all 217 AI-generated vs. human-reviewed helper pairs using an external rater (GPT-4o-mini via OpenRouter). The rater assigned 0–10 scores on four dimensions aligned with our quality rubric: accuracy, clarity, completeness, and reliability. Human-reviewed passages outperform AI-generated passages by +1.38 to +1.87 points per metric and by +1.72 points on the overall average (Table 6), confirming that expert edits measurably increase helper quality beyond text length controls.
Table 6: LLM-based quality ratings (0–10) for 217 helper pairs. Scores are means across questions.
Metric AI helper Human helper $\boldsymbol{\Delta}$
Accuracy 7.43 9.25 +1.82
Clarity 7.53 8.91 +1.38
Completeness 7.40 9.20 +1.81
Reliability 7.30 9.17 +1.87
Overall mean 7.41 9.13 +1.72
3.5.3 Differential Performance Under Token-Matched Conditions
Table 7 presents a direct comparison of performance gains ($\Delta$) between token-matched AI-generated and expert-reviewed supporting materials across all 74 evaluated models. Despite identical token counts, substantial performance differences emerge, revealing the effect of content quality independent of text volume.
Most models show higher accuracy with expert-reviewed materials, confirming that qualitative improvements—such as enhanced conceptual clarity, precise terminology, and improved contextual alignment—produce measurable performance gains even under identical token counts. The strongest relative improvements occur in small models (<7B parameters), which benefit most from the additional clarity and precision introduced by expert curation. For instance, LLaVA-v1.6-mistral-7B improves from +5.54% to +11.07% (+5.53% differential), and mistralai/mistral-7b-instruct-v0.1 exhibits the largest quality-driven transformation (+8.41%), shifting from performance degradation (–5.47%) with AI-generated text to a measurable gain (+2.94%) with expert-reviewed material, under identical token budgets.
Mid-sized models (10–70B) also show clear but smaller improvements, exemplified by LLaVA-v1.6-34B (+5.99% to +11.52%, +5.53% differential). These models possess enough reasoning capacity to leverage higher-quality information but already contain partial crystallographic knowledge, so their absolute gains are moderate.
Conversely, large models (>70B), such as GPT-4.5-preview, O3-mini, and O1, exhibit minimal or slightly negative differences, reflecting a saturation of internal knowledge representations. When most relevant crystallographic concepts are already encoded within the model parameters, additional context contributes little new information and can occasionally disrupt attention alignment or introduce representational interference.
Overall, these results confirm that content quality, not text length, governs open-book performance. Expert review yields the greatest relative benefits for small and mid-sized models, while large models approach a ceiling where additional context produces diminishing or even adverse effects.
Figure 10 summarizes the overall impact of supporting material quality across model size categories. Small models (<7B parameters) achieve the largest relative improvement from expert-reviewed materials, increasing their accuracy by +8.52 percentage points compared to +6.18% with AI-generated content. Mid-sized models (7B–70B) also benefit noticeably, showing a +4.44% gain versus +2.71% with AI-generated materials. In contrast, large models (>70B) display minimal or slightly negative changes with either material type, indicating that their extensive internal knowledge already limits the marginal benefit of added context.
This trend can be explained by differences in knowledge representation and context utilization capacity across model scales. Smaller models lack sufficient internal crystallographic knowledge, so expert-reviewed passages provide missing conceptual structure and precise terminology that directly enhance reasoning accuracy. Mid-sized models already encode partial domain knowledge but still rely on external information to fill gaps, making them responsive to high-quality supporting text. In contrast, very large models possess extensive parametric knowledge, and adding redundant or overlapping context can introduce interference effects, where new information competes with their established internal representations. As a result, they gain little additional benefit and may even experience slight degradation when the external input does not perfectly align with their internal reasoning pathways.
Overall, these findings indicate that while expert curation improves performance across all model scales, the relative gains are most pronounced for small and mid-sized models that can still meaningfully integrate external guidance, whereas large models operate near their knowledge saturation limit.
3.5.3.1 Practical significance of Fig. 10.
Improving average accuracy from 72.5% to 79.5% (Fig. 10) is a 7.0 percentage-point gain that lowers the error rate from 27.5% to 20.5%—about a 25% reduction in wrong answers, or roughly 15 additional correct answers on our 217-question benchmark. Moving from 3 wrong answers in 10 to 2 can be noticeable to users because it reduces re-checking and re-asking. However, 79.5% still means 1 in 5 answers is incorrect, so high-stakes crystallography (e.g., indexing, space-group interpretation, experimental planning) should remain human-verified. The main benefit is cost-effective assistance: mid-sized models plus expert-reviewed hints approach large-model accuracy at lower compute cost.
3.5.4 Quantity-Driven Effects: Evidence from Token Budget Constraints
To further disentangle quality from quantity effects, Section 3.7 presents complementary evidence that information quantity alone can degrade performance. Our token budget analysis (Figure 11) reveals that llamat-3-chat’s open-book accuracy dropped from 36.87% at 256 tokens to 26.73% at 2048+ tokens—a 10 percentage point degradation—when provided with longer supporting materials. This inverse relationship demonstrates severe token starvation: the model exhausts its computational budget on input processing, leaving insufficient capacity for answer generation. The supporting materials, rather than aiding reasoning, become a computational burden that crowds out response formulation.
This finding establishes a critical control, if performance gains stemmed primarily from text volume rather than content quality, we would not observe degradation with increased token budgets in resource-constrained models. The token starvation phenomenon proves that “more text” does not guarantee better outcomes and, in fact, can be counterproductive when models lack the capacity to efficiently process additional context.
3.5.5 Interpretation: Quality, Not Quantity, Drives Improvements
The combination of token-matched comparisons (identical length, differential outcomes) and token budget analysis (increased length, degraded outcomes) provides convergent evidence that content quality—specifically, accuracy, relevance, and pedagogical clarity—drives the observed performance improvements, not mere text volume. Expert-reviewed materials successfully addressed the limitations identified in AI-generated content: technical inaccuracies, insufficient depth, tangential information, and suboptimal pedagogical framing. These qualitative enhancements enable mid-capacity models to better integrate external knowledge, bridging the gap between their processing capability and incomplete domain expertise.
Critically, the token-matched design ensures that expert improvements cannot be attributed to providing "more information" in terms of token count. Instead, expert review enhances information density, precision, and alignment with question-specific reasoning needs—dimensions of quality orthogonal to text length. This ablation establishes that the OPENXRD framework’s performance gains reflect genuine knowledge enhancement rather than computational artifacts of longer context windows.